OpenStudy (tootsi123):

Christine went shopping and bought each of her eight nephews a gift., either a video costing $14.95 or a cd costing $16.88. She spent $131.18 on the gifts. How many videos and how many cds did she buy?

OpenStudy (mathstudent55):

Let v = number of videos, and let c = number of cd's. How many total videos and cd's did she buy?

OpenStudy (tootsi123):

I have no clue..

OpenStudy (mathstudent55):

You need to read the problem before you can answer it. It says how many.

OpenStudy (mathstudent55):

"Christine went shopping and bought each of her eight nephews a gift..."

OpenStudy (tootsi123):

It says how much it costs?

OpenStudy (mathstudent55):

We'll deal with the price later. Now we are just dealing with the numbers of videos and cd's.

OpenStudy (tootsi123):

Okay so she bought 8 items, that still doesnt help me ?

OpenStudy (mathstudent55):

You are correct. It is 8. It does help.

OpenStudy (mathstudent55):

v is the number of videos. c is the number of cd's. We don;t know yet how many videos and how many cd's she bought, but we do know the total number she bought was 8. That means we can write this equation: v + c = 8 Ok so far?

OpenStudy (tootsi123):

Yes.

OpenStudy (mathstudent55):

Now we start thinking of prices and the amount of money spent. If one video costs $14.95, how much do v videos cost?

OpenStudy (tootsi123):

14.95

OpenStudy (tootsi123):

14.95+16.88=131.18??

OpenStudy (mathstudent55):

No. Just follow along. 14.95 is the price of 1 video. v is the number of videos bought. For example, let's say she bought 5 videos. 1 video costs 14.95. What operation would you do to find how much 5 videos cost?

OpenStudy (tootsi123):

14.95+5=8?

OpenStudy (mathstudent55):

No. If 1 video costs 14.95, then 1 video costs 1 * 14.95 2 videos cost 2 * 14.95 3 videos cost 3 * 14.95 4 videos cost 4 * 14.95 etc. You multiply the number of videos by how much each one costs to find the price of all the videos. That means 5 videos cost 5 * 14.95 Since we don't know yet how many videos were bought, we are calling that number v. Then, the price paid for all the videos bought is v (the number of videos bought) times $14.95. 2 videos cost 2 * 14.95 v videos cost v * 14.95, or 14.95v

OpenStudy (tootsi123):

Okay

OpenStudy (mathstudent55):

Each video costs 14.95. v videos cost 14.95v Now we do the same with the cd's Each cd costs $16.88 Then c number of cd's cost 16.88c

OpenStudy (tootsi123):

ok

OpenStudy (mathstudent55):

The total amount spent, then, is \(\sf \color{red}{the ~amount ~spent ~on ~videos}~ plus ~\color{green}{the ~amount ~spent ~on ~cd's}\) \(\color{red}{~~~~~~~~~~~~~~~~~14.95v} ~~~~~~~~~~~~~~~~~~~~+ \color{green}{~~~~~~~~~~~~~16.88c}\)

OpenStudy (mathstudent55):

Without the colors, the amount spent is 14.95v + 16.88c The problem tells us that the total amount spent was $131.18, that means that 14.95v + 16.88c must equal 131.88 That gives us the second equation: 14.95v + 16.88c = 131.88

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (tootsi123):

yea

OpenStudy (mathstudent55):

Now we need to use the two equations below as a system of equations and solve them to find v and c. \(v + c = 8\) \(14.95v + 16.88c = 131.18\)

OpenStudy (mathstudent55):

Have you learned the substitution method of solving systems of equations?

OpenStudy (tootsi123):

yea

OpenStudy (tootsi123):

Thanks. I got it.

OpenStudy (mathstudent55):

Solve the first equation for v: c = 8 - v Substitute 8 - v in for c in the second equation: 14.95v + 16.88(8 - v) = 131.18 14.95v + 135.04 - 16.88v = 131.18 -1.93v = 3.86 v = 2 Now substitute v = 2 in v + c = 8 to find c.